Session: May 1998 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Maccaferri Andrea at Sun May 31 12:20:38 METDST 1998) Success
good weather

Torun (K. Borkowski at Sun May 31 13:45:57 METDST 1998) Failure
No observations due to a failure in the telescope elevation drive.

Onsala (A. Polatidis at Sun May 31 14:05:01 METDST 1998) Success
No known problems. Very good weather.

Westerbork (Tony Foley at Sun May 31 20:56:35 METDST 1998) Failure
Only final scan recorded (07:00-07:22) because of tape running off the reel on previous experiment and operator error.

Jodrell (Alastair gunn at Mon Jun 1 15:43:47 METDST 1998) Success
No known problems.

Westerbork (dgraham at Tue Jun 2 18:39:01 METDST 1998) Minor failures

Calibration (Jean\-Francois Desmurs at Mon Jun 8 13:26:00 METDST 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the may98 directory under the name v030z.antab. Read the comments in the file to get further information.

Noto (C.Trigilio at Mon Jun 8 17:24:15 METDST 1998) Success